Aspire Faculty ID #16294 · Topic: NIMCET 2010 · Just now
NIMCET 2010

India plays 4 matches. Probabilities of scoring $0,1,2$ points: $0.45,\ 0.05,\ 0.50$. Find probability of at least $7$ points.

Solution

$P(8)=0.5^4=0.0625$ 
$P(7)=4 \cdot 0.05 \cdot 0.5^3 = 0.025$ 
Total $= 0.0625 + 0.025 = 0.0875$

Previous 10 Questions — NIMCET 2010

Nearest first

Next 10 Questions — NIMCET 2010

Ascending by ID
Ask Your Question or Put Your Review.

loading...